OpenStudy (anonymous):

Evaluate for x = 2 and y = 3 X^2Y^-3/X^-1Y 8/81 2/9 1/648

OpenStudy (anonymous):

Evaluate for x = 2 and y = 3 X^2Y^-3/X^-1Y

OpenStudy (anonymous):

8/81 2/9 1/648

OpenStudy (anonymous):

Is it the expression?\[\frac{ x^{2}y^{-3} }{ x^{-1} y}\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Is it the expression? YES

OpenStudy (anonymous):

I NEED HELP Zekarias

OpenStudy (anonymous):

\[\frac{ x^{2}y^{-3} }{ x^{-1} y}=x^{3}y^{-4}\]\[=2^{3}3^{-4}\]\[=8{\times}81^{-1}\]\[=\frac{ 8 }{ 81 }\]

OpenStudy (anonymous):

Did you get ?

OpenStudy (anonymous):

THANKS